OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points. (1, 7), (10, 1)

OpenStudy (pixiedust1):

Okay! So the formula for slope is \[\frac{ y _{2} -y _{1}}{ x _{2} -x _{1}}\]

OpenStudy (pixiedust1):

Do you know what the next step is?

OpenStudy (pixiedust1):

I see that you closed the question. For all the google users out there, here is the rest of the problem. y2=1 Why? Because 1 is the second number of the second pair of points. y1= 10 Why? Because ts the first number of the second pair of points. 1-10=-9 x2=7 Why? Because 7 is the second number of the first pair of points. x1=1. Why? Because 1 is the first number of the first pair of points. 7-1=6 Now, we have \[\frac{ -9 }{ 6}\]

OpenStudy (pixiedust1):

Therefore, the slope is -1.5